Use the identity (x + a) (x + b) = x+ (a + b) x + ab to find the following products.
(x + 3) (x + 7)
 
Please solve this fast

Grade:8

2 Answers

Ayush 8A 7239
2 years ago
Hi mate, here is the answer to the question
 
Using identity (x + a) (x + b) = x2 + (a + b) x + ab
(x+3) (x+7) = x2 + (3+7)x + 3×7
                 = x2 + 10x + 21
 
Hope it helps

 

Ayush 8A 7239
2 years ago
Using identity (x + a) (x + b) = x2 + (a + b) x + ab
(x+3) (x+7) = x2 + (3+7)x + 3×7
                 = x2 + 10x + 21
 
Hope it helped you.
Thanks